MONOBRIDGED Si_{2}H_{4}

\underline{M. C. McCARTHY} and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; LEVENT SARI and HENRY F. SCHAEFER III, Center for Computational Quantum Chemistry, University of Georgia, Athens, Georgia 30602-2525; ZHENHONG YU, Department of Chemistry, Harvard University, Cambridge, MA 02138.

The rotational spectrum of a novel monobridged isomer of Si_{2}H_{4}, denoted here as H_{2}SiHSiH, has been detected by means of Fourier transform microwave spectroscopy in a supersonic molecular beam through the discharge products of silane. On the basis of high-level coupled cluster theory, this isomer is calculated to lie only 7 kcal/mol above disilene (H_{2}SiSiH_{2}), the most stable isomeric arrangement of Si_{2}H_{4}, and to be fairly polar, with a calculated dipole moment of μ = 1.14 D. The rotational spectrum of H_{2}SiHSiH exhibits closely-spaced line doubling, characteristic of a molecule undergoing high-frequency inversion motion; at the CCSD(T) level of theory the barrier height to inversion is calculated to be 10-11 kcal/mol. Because of its favorable energetics and high polarity with respect to either disilene or silylsilylene, monobridged Si_{2}H_{4} is a good candidate for radioastronomical detection in sources such as IRC+10216 where other large silicon-bearing molecules have already been discovered.